A 21 inch square TV is on sale at the local electronics store. If 21 inches is the measure of the diagonal of the screen, use the Pythagorean theorem to find the length of the side of the screen.
Answer by ewatrrr(24785)

 
Hi,
A 21 inch square TV is on sale at the local electronics store.
Let x represent the length of the side of the screen
Using the Pythagorean theorem:
x^2 + x^2 = 21^2
2x^2 = 441
x^2 = 220.5
x = sqrt(220.5) = 14.85in